# S2ST-Omni: An Efficient and Scalable Speech-to-Speech Translation Framework via Seamlessly Speech-Text Alignment and Streaming Speech Decoder

### v2

@@ Abstract
Multilingual speech-to-speech translation (S2ST) aims to directly convert spoken utterances from multiple source languages into fluent and intelligible speech in a target language. Despite recent progress, several critical challenges persist: 1) achieving high-quality and low-latency S2ST remains a significant obstacle; 2) most existing S2ST methods rely heavily on large-scale parallel speech corpora, which are difficult and resource-intensive to obtain. To tackle these challenges, we introduce S2ST-Omni, a novel, efficient, and scalable framework tailored for multilingual speech-to-speech translation. To enable high-quality S2TT while mitigating reliance on large-scale parallel speech corpora, we leverage powerful pretrained models: Whisper for robust audio understanding and Qwen 3.0 for advanced text comprehension. A lightweight speech adapter is introduced to bridge the modality gap between speech and text representations, facilitating effective utilization of pretrained multimodal knowledge. To ensure both translation accuracy and real-time responsiveness, we adopt a streaming speech decoder in the TTS stage, which generates the target speech in an autoregressive manner. Extensive experiments conducted on the CVSS benchmark demonstrate that S2ST-Omni consistently surpasses several state-of-the-art S2ST baselines in translation quality, highlighting its effectiveness and superiority.
@@ 3.3 TTS Stage: Streaming Speech Decoder
ASR-BLEU $\uparrow$ ASR-BLEU $\uparrow$ ASR-BLEU $\uparrow$
S2UT [32] 22.23 2.99 18.53
Translatotron 2 [17] 26.12 16.92 22.98
DASpeech [14] 25.03 16.14 21.37
UnitY [18] 26.90 16.36 24.06
ComSpeech [12] 28.15 18.16 24.80
StreamSpeech [15] 28.45 20.93 27.25
S2ST-Omni 67.45 71.20 68.28
@@ 4.1.2 Implementation Details
Regarding the evaluating metric, we employ the commonly used ASR-BLEU to assess performance. In practice, the translated speech is first transcribed into its corresponding transcript using a pre-trained ASR model4, after which the BLEU score [36] is computed using the SacreBLEU toolkit5 to assess translation quality and statistical significance.
@@ 4.2 Results and Analysis
To validate the effectiveness of the proposed approach, we compare the translation performance of S2ST-Omni against six state-of-the-art S2ST systems. A comprehensive summary of the results is provided in Table 1.
As shown in the table, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with gains exceeding 38 BLEU points in certain cases. These results highlight the strong generalization capability of the proposed framework across linguistically diverse settings. Specifically, in the French $\rightarrow$ English direction, S2ST-Omni achieves an ASR-BLEU score of 67.45, markedly surpassing the strongest baseline (StreamSpeech, 28.45). This substantial improvement underscores the model’s robustness in handling phonetic variability and prosodic features characteristic of French. For German $\rightarrow$ English, the model attains a remarkable 71.20 ASR-BLEU score—the highest among all language pairs—demonstrating its effectiveness in managing syntactic complexity and compound structures inherent to German. In the Spanish $\rightarrow$ English setting, S2ST-Omni achieves a strong 68.28 ASR-BLEU score, again outperforming all baselines by a wide margin. This result reflects the model’s ability to capture the rich morphological patterns of Spanish and generate fluent, accurate translations.
The outstanding performance of S2ST-Omni can be attributed to several key architectural innovations:
  • Leveraging large-scale pre-trained models: By integrating Whisper’s robust acoustic representation learning with the powerful semantic modeling capabilities of Qwen 3.0, S2ST-Omni effectively exploits the rich knowledge embedded in large-scale pre-training. This substantially reduces the reliance on extensive parallel S2ST data, which is often scarce and costly to obtain.
  • Seamless speech-text alignment: The introduction of lightweight speech adapters enables the effective mapping of Whisper’s speech representations into the language space of Qwen 3.0. This design bridges the modality gap between speech and text, facilitates efficient knowledge transfer, and enhances alignment quality for downstream translation tasks.
  • Streaming speech decoder for TTS: A streaming speech decoder is employed during the text-to-speech synthesis stage, achieving real-time, low-latency voice generation without compromising translation quality—an essential requirement for practical deployment.
@@ 5 CONCLUSIONS
In this study, we introduce S2ST-Omni, a unified and efficient framework for multilingual speech-to-speech translation. By decomposing the S2ST task into speech-to-text translation and text-to-speech synthesis, and integrating large-scale pre-trained models such as Whisper and Qwen 3.0 with a lightweight speech adapter, S2ST-Omni achieves high-quality translation while reducing reliance on parallel speech data. Furthermore, the adoption of a streaming speech decoder enables low-latency speech generation suitable for real-time applications. Extensive experiments on the CVSS benchmark confirm that S2ST-Omni achieves better translation quality than existing state-of-the-art systems, demonstrating its effectiveness and superiority in practical multilingual S2ST applications.

### v3

# S2ST-Omni: An Efficient and Scalable Multilingual Speech-to-Speech Translation Framework via Seamless Speech-Text Alignment and Streaming Speech Generation

@@ Abstract
Multilingual speech-to-speech translation (S2ST) aims to directly convert spoken utterances from multiple source languages into fluent and intelligible speech in a target language. Despite recent progress, several critical challenges persist: 1) achieving high-quality S2ST remains a significant obstacle; 2) most existing S2ST methods rely heavily on large-scale parallel speech corpora, which are difficult and resource-intensive to obtain. To tackle these challenges, we introduce S2ST-Omni, a novel, efficient, and scalable framework tailored for multilingual speech-to-speech translation. Specifically, we decompose S2ST into speech-to-text translation (S2TT) and text-to-speech synthesis (TTS). To enable high-quality S2TT while mitigating reliance on large-scale parallel speech corpora, we leverage powerful pretrained models: Whisper for robust audio understanding and Qwen 3.0 for advanced text comprehension. A lightweight speech adapter is introduced to bridge the modality gap between speech and text representations, facilitating effective utilization of pretrained multimodal knowledge. To ensure both translation accuracy and real-time responsiveness, we adopt a streaming speech generation model in the TTS stage, which generates the target speech in an autoregressive manner. Extensive experiments conducted on the CVSS benchmark demonstrate that S2ST-Omni consistently surpasses several state-of-the-art S2ST baselines in translation quality, highlighting its effectiveness and superiority.
@@ 3.3 TTS Stage: Streaming Speech Decoder
S2UT [32] 22.23 2.99 18.53 14.58
Translatotron 2 [17] 26.12 16.92 22.98 22.01
DASpeech [14] 25.03 16.14 21.37 20.85
UnitY [18] 26.90 16.36 24.06 22.44
ComSpeech [12] 28.15 18.16 24.80 23.70
StreamSpeech [15] 28.45 20.93 27.25 25.54
S2ST-Omni 67.45 71.20 68.28 68.98
@@ 4.1.2 Implementation Details
Regarding evaluation, we adopt two widely used metrics: ASR-BLEU and BLASER 2.0, to comprehensively assess the performance of all comparative methods. Specifically, for ASR-BLEU, the translated speech is first transcribed into text using a pretrained ASR model4. The resulting transcripts are then evaluated using the BLEU score [36], calculated with the SacreBLEU toolkit5, to quantify translation quality and statistical significance. For BLASER, we follow [15], employing the BLASER 2.0 toolkit6 to measure cross-lingual semantic similarity between source and translated utterances.
@@ 4.1.3 Baselines
Table 2: Offline S2ST performance of StreamSpeech, evaluated with BLASER 2.0.
[…]
BLASER 2.0 (Unsupervised)
UnitY 0.4467 0.4473 0.5090 0.5116
StreamSpeech 0.4486 0.4491 0.5155 0.5178
S2ST-Omni 0.6523 0.7138 0.7262 0.6974
BLASER 2.0 (QE)
UnitY 3.1674 3.1772 3.3020 3.3278
StreamSpeech 3.1779 3.1872 3.3442 3.3669
S2ST-Omni 3.6475 3.9284 3.9648 3.8469
BLASER 2.0 (Ref)
UnitY 3.1744 3.1965 3.2213 3.2638
StreamSpeech 3.1989 3.2200 3.3146 3.3525
S2ST-Omni 3.8925 4.0897 4.0634 4.0152
@@ 4.2 Results and Analysis
To validate the effectiveness of the proposed approach, we compare the translation performance of S2ST-Omni against six state-of-the-art S2ST systems. A comprehensive summary of the results is provided in the Table 1 and Table 2.
As shown in Table 1, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with ASR-BLEU improvements exceeding in certain cases. These results highlight the strong generalization capability of the proposed framework across linguistically diverse settings. Specifically, in the French $\rightarrow$ English direction, S2ST-Omni achieves an ASR-BLEU score of 67.45, markedly surpassing the strongest baseline (StreamSpeech, 28.45). This substantial improvement underscores the model’s robustness in handling phonetic variability and prosodic features characteristic of French. For German $\rightarrow$ English, the model attains a remarkable 71.20 ASR-BLEU score—the highest among all language pairs—demonstrating its effectiveness in managing syntactic complexity and compound structures inherent to German. In the Spanish $\rightarrow$ English setting, S2ST-Omni achieves a strong 68.28 ASR-BLEU score, again outperforming all baselines by a wide margin. This result reflects the model’s ability to capture the rich morphological patterns of Spanish and generate fluent, accurate translations.
Furthermore, to directly evaluate the quality of the converted speech, we also conduct BLASER comparisons which contain Unsupervised (score from 0 to 1), QE (score from 1 to 5),and Ref (score from 1 to 5) modes. As illustrated in Table 2, we can easily observe that the proposed S2ST-Omni system consistently achieves the best performance across all evaluation modes. In particular, S2ST-Omni achieves 34.7%, 14.3% and 19.8% relative BLASER improvements over the secondary StreamSpeech in terms of Unsupervised, QE, and Ref modes, highlighting the effectiveness and superiority of the proposed approach.
@@ 5 CONCLUSIONS
In this study, we introduce S2ST-Omni, a unified and efficient framework for multilingual speech-to-speech translation. By decomposing the S2ST task into speech-to-text translation and text-to-speech synthesis, and integrating large-scale pre-trained models such as Whisper and Qwen 3.0 with a lightweight speech adapter, S2ST-Omni achieves high-quality translation while reducing reliance on parallel speech data. Furthermore, the adoption of a streaming speech generation model enables low-latency speech generation suitable for real-time applications. Extensive experiments on the CVSS benchmark confirm that S2ST-Omni achieves better translation quality than existing state-of-the-art systems, demonstrating its effectiveness and superiority.